User name LeRose
Log entry time 16:25:26 on December 22,1999
Entry number 31829
The following procedure was used for all spectrometer settings in E89-044.
Electron Q2/3's were run up to 1600 Amps and then down to the desired current.
Hadron Q2/3's were run up to 1400 Amps and then down to the desired current.
No cycling on Q1's or Dipoles (both arms) (according to Arun)